### v1

# Multi-Source Position and Direction-of-Arrival Estimation Based on Euclidean Distance Matrices

Abstract
A popular method to estimate the positions or directions-of-arrival (DOAs) of multiple sound sources using an array of microphones is based on steered-response power (SRP) beamforming. For a three-dimensional scenario, SRP-based methods need to jointly optimize three continuous variables for position estimation or two continuous variables for DOA estimation. This can be computationally expensive, especially when high localization accuracy is desired. In this paper, we propose novel methods for multi-source position and DOA estimation by exploiting properties of Euclidean distance matrices (EDMs) and their respective Gram matrices. All methods require estimated time-differences of arrival (TDOAs) between the microphones. In the proposed multi-source position estimation method only a single continuous variable, representing the distance between each source and a reference microphone, needs to be optimized. For each source, the optimal continuous distance variable and set of candidate TDOA estimates are determined by minimizing a cost function that is defined using the eigenvalues of the Gram matrix. The estimated relative source positions are then mapped to estimated absolute source positions by solving an orthogonal Procrustes problem for each source. The proposed multi-source DOA estimation method entirely eliminates the need for continuous variable optimization by defining a relative coordinate system per source such that one of its coordinate axes is aligned with the respective source DOA. The optimal set of candidate TDOA estimates is determined by minimizing a cost function that is defined using the eigenvalues of a rank-reduced Gram matrix. The computational cost of the proposed EDM-based methods is significantly reduced compared to the SRP-based methods, and for EDM-based DOA estimation the need for continuous variable optimization is even entirely eliminated. For two sources in a noisy and reverberant environment, experimental results for different source and microphone configurations show that the proposed EDM-based method consistently outperforms the SRP-based method in terms of position and DOA estimation accuracy.
I Introduction
Estimating the locations, i.e., positions or directions-of-arrival (DOAs), of multiple speech sources in a noisy and reverberant environment is important for several applications, such as source tracking, speech enhancement, and speaker extraction [1, 2, 3, 4]. In this paper, we will consider both compact microphone arrays with relatively small inter-microphone distances as well as spatially distributed microphones with large inter-microphone distances. For compact microphone arrays the sources will be assumed to be in the far field, such that we will focus on estimating the DOA (azimuth and elevation) of each source, whereas for spatially distributed microphones we will focus on estimating the three-dimensional position of each source.
Existing model-based methods for source localization fall into two main categories [1, 2]: one-step methods such as steered-response power methods [5, 6, 7] or the subspace-based multiple signal classification (MUSIC) method [8], and two-step methods, which rely on the prior estimation of variables such as time-differences of arrival (TDOAs) between microphones [9]. In addition, recently several learning-based methods [10, 11, 12, 13, 14, 15, 16] for source localization have been proposed, showing promising results. However, most learning-based methods are trained for a specific array geometry and do not support source localization with arbitrary array geometries. In this paper, we only consider model-based localization methods, which provide the flexibility to use different array geometries without the need to retrain a neural network. A popular model-based method for single-source DOA or position estimation is based on the steered-response power with phase transform (SRP-PHAT) functional [6, 17, 18, 19, 20, 21, 22, 23]. While DOA estimation requires jointly optimizing the SRP-PHAT functional in up to two continuous variables (azimuth and elevation), position estimation requires jointly optimizing the SRP-PHAT functional in up to three continuous variables (x, y, and z coordinates). To perform accurate source localization, the SRP-PHAT functional needs to be evaluated on a discrete multi-dimensional grid with a high enough resolution in each variable resulting in a high computational complexity [21, 24]. To avoid the multi-dimensional search of SRP-based methods, in [25] we proposed a method to estimate the three-dimensional position of a single source based on Euclidean distance matrices. This method minimizes a cost function defined using the eigenvalues of a Gram matrix associated with an EDM containing the inter-microphone distances and the source-microphone distances. Using estimated TDOAs, this problem can be reformulated as an optimization in only a single distance variable, where the solution represents the distance between the source and the reference microphone. Several model-based approaches have also been proposed for multi-source DOA and position estimation, e.g., based on determining multiple peaks in the SRP-PHAT functional or MUSIC spectrum [8, 7, 26, 27, 28, 29]. It should however be noted that most of these approaches focus on multi-source DOA estimation and not on multi-source position estimation. Moreover, many approaches rely on data association/clustering of acoustic features [26, 30, 31] and are quite sensitive to reverberation.
In this paper, we first extend the EDM-based single-source position estimation method in [25] to multi-source position estimation, assuming the number of sources $S$ to be known. For each combination of candidate TDOA estimates, the optimal distance variable is determined by minimizing the EDM-based cost function using a one-dimensional exhaustive search. Assuming that each source position corresponds to a unique combination of TDOAs, the optimal set of candidate TDOA estimates and corresponding source distances are determined, thereby also solving the association of candidate TDOA estimates to sources. The estimated relative position of each source to the microphone array can then be determined from the Gram matrix with the estimated source distance. The estimated relative positions are then mapped to estimated absolute source positions by solving an orthogonal Procrustes problem for each source. As the second contribution of this paper, we propose an EDM-based method to estimate the DOAs of multiple sources relative to a compact microphone array, which does not require continuous variable optimization. We construct a rank-reduced Gram matrix by subtracting a rank-1 matrix based on the estimated TDOAs from the Gram matrix associated with the EDM containing the microphone distances. We define a cost function using the eigenvalues of this rank-reduced Gram matrix and, similarly as for multi-source position estimation, determine the optimal set of candidate TDOA estimates by considering the $S$ smallest values of this cost function. The DOA of each source is then estimated from the mapping between the absolute coordinate system and the relative coordinate system, which depends on the combinations of TDOA estimates corresponding to each source.
Two sets of experiments are conducted for several simulated scenarios with two sources and six microphones in a room with mild background noise and reverberation. The results of the first experiment considering spatially distributed microphones show that the proposed EDM-based method consistently outperforms the SRP-based method in estimating the of the sources for all considered positions. Similarly, the results of the second experiment considering compact microphone arrays show that the proposed EDM-based method also outperforms the SRP-based method in estimating the DOAs of the sources for all considered distances between the sources and the microphone array. Since the EDM-based methods reduce or eliminate the number of continuous variables that need to be optimized, the runtime can be considerably reduced compared to the SRP-based methods (about 380 times for position estimation and about 25 times for DOA estimation).
This paper is organized as follows. In Section II, we introduce the acoustic scenario and the theoretical background and properties of EDMs. After reviewing the EDM-based position estimation method for a single source, we extend this method for multi-source position estimation in Section III. In Section IV we propose EDM-based methods for single-source and multi-source DOA estimation considering far-field sources and compact microphone arrays. After discussing the baseline SRP-based method in Section V, we compare the performance of the EDM-based and SRP-based methods for multi-source position and DOA estimation in Section VI.
II Theoretical Background
We consider a noisy and reverberant acoustic environment with $S$ static sources, where $\mathbf{p}_{s}\in\mathbb{R}^{P}$ denotes the position of the $s$-th source in the absolute coordinate system with $P$-dimensional canonical basis vectors $\mathbf{e}_{1},\dots,\mathbf{e}_{P}$ ($1\leq P\leq 3$). We assume the number of sources $S$ to be known. We consider an array with $M$ microphones whose positions are assumed to be known, with $M>P$, where $\mathbf{m}_{m}\in\mathbb{R}^{P}$ denotes the position of the $m$-th microphone. The $P\times M$-dimensional microphone positions matrix is defined as
$\mathbf{M}\;=\;[\mathbf{m}_{1},\dots,\mathbf{m}_{M}]\;.$ (1)
The origin of the coordinate system is defined at the centroid of the microphone positions, i.e., $\frac{1}{M}\mathbf{M}\mathbf{1}_{M}=\mathbf{0}_{P}$, with $\mathbf{1}_{M}$ denoting an $M$-dimensional vector of ones and $\mathbf{0}_{P}$ denoting a $P$-dimensional vector of zeros. The source and microphone positions are exemplified in Fig. 1.
[…]
Figure 1: Exemplary acoustic scenario with two sources at positions $\mathbf{p}_{1}$ and $\mathbf{p}_{2}$ and a microphone array at positions $\mathbf{m}_{1},\dots,\mathbf{m}_{M}$, with the first microphone defined as the reference microphone.
The $P\times(M+1)$-dimensional microphones and source positions matrix for the $s$-th source is defined as $\mathbf{P}_{s}=[\mathbf{M},\;\mathbf{p}_{s}]$. The $M\times M$-dimensional Gram matrix of the microphones $\mathbf{G}_{MM}$ and the $(M+1)\times(M+1)$-dimensional Gram matrix for the $s$-th source $\mathbf{G}_{s}$ are defined as
$\mathbf{G}_{MM}\;=\;\mathbf{M}^{\textrm{T}}\mathbf{M}\;,$ (2)
$\mathbf{G}_{s}\;=\;\mathbf{P}_{s}^{\textrm{T}}\mathbf{P}_{s}\;,$ (3)
where $\{\cdot\}^{\textrm{T}}$ denotes the transpose operator. An important property of both Gram matrices, which will be used throughout this paper, is that their rank is at most $P$. When considering another orthonormal basis, relative to the original canonical basis (obtained by rotating and/or reflecting the original basis), the relative microphone positions matrix $\mathbf{M}_{\textrm{r}}$ and the relative microphones and source positions matrix for the $s$-th source $\mathbf{P}_{s}$ are given by
$\mathbf{M}_{\textrm{r}}\;=\;\mathbf{R}^{\textrm{T}}\mathbf{M}\;,\;$ (4)
$\mathbf{P}_{\text{r}s}\;=\;\mathbf{R}^{\textrm{T}}\mathbf{P}_{s}\;=\;[\mathbf{M}_{\textrm{r}}\,,\;\mathbf{R}^{\textrm{T}}\mathbf{p}_{s}]\;,$ (5)
with $\mathbf{R}$ a $P\times P$-dimensional orthogonal matrix, for which $\mathbf{R}^{\textrm{T}}=\mathbf{R}^{-1}$.
It is important to realize that the Gram matrices in (2) and (3) are unaffected by a basis rotation and/or reflection, i.e.,
$\mathbf{G}_{MM}\;=\;\mathbf{M}_{\textrm{r}}^{\textrm{T}}\mathbf{M}_{\textrm{r}}\;,$ (6)
$\mathbf{G}_{s}\;=\;\mathbf{P}_{\textrm{r}s}^{\textrm{T}}\mathbf{P}_{\textrm{r}s}\;.$ (7)
[…]
where $\nu$ denotes the speed of sound. Without loss of generality, we define the first microphone as the reference microphone. From (8), it can be directly seen that the distance between the $s$-th source and the $m$-th microphone can be written in terms of the distance between the $s$-th source and the reference microphone as
$d_{ms}\;=\;d_{1s}\;+\;\nu\tau_{m}(\mathbf{p}_{s})\;,$ (9)
with $\tau_{m}(\mathbf{p}_{s})$ denoting the TDOA between the $m$-th microphone and the reference microphone. The $(M+1)\times(M+1)$-dimensional EDM $\mathbf{D}_{s}$ for the $s$-th source is defined as
${\mathbf{D}}_{s}\;=\;\begin{bmatrix}\begin{array}[]{c|c}\mathbf{D}_{MM}&\mathbf{d}_{s}\\ \hline\cr\mathbf{d}_{s}^{\textrm{T}}&0\end{array}\end{bmatrix}\;$, (10)
where the submatrix $\mathbf{D}_{MM}=\left[D^{2}_{ij}\right]$ contains the squared inter-microphone distances and the vector $\mathbf{d}_{s}=\left[d^{2}_{1s},d^{2}_{2s},\dots,d^{2}_{Ms}\right]^{\textrm{T}}$ contains the squared distances between the $s$-th source and the microphones. In [32, 33], it was shown that the Gram matrix of the microphones ${\mathbf{G}}_{MM}$ in (2) could be written in terms of the EDM $\mathbf{D}_{MM}$ as
$\mathbf{G}_{MM}\;=\;-\frac{1}{2}\,(\mathbf{I}_{M}-\mathbf{1}_{M}\mathbf{a}_{M}^{\textrm{T}})\,{\mathbf{D}_{MM}}\,(\mathbf{I}_{M}-\mathbf{a}_{M}\mathbf{1}_{M}^{\textrm{T}})\;,$ (11)
where $\mathbf{I}_{M}$ denotes the $M\times M$-dimensional identity matrix and $\mathbf{a}_{M}$ denotes the so-called centering vector. To ensure that the centroid of the relative microphone positions $\mathbf{M}_{\textrm{r}}$ is at the origin, we define the centering vector for $\mathbf{G}_{MM}$ as $\mathbf{a}_{M}=\frac{1}{M}\mathbf{1}_{M}$. This important property will be used in the EDM-based DOA estimation in Section IV.
For the EDM-based position estimation, we extend (11) to include the distances between the $s$-th source and the microphones, i.e., for Similarly as in (11), the Gram matrix for the $s$-th source $\mathbf{G}_{s}$ can be written in terms of the EDM $\mathbf{D}_{s}$ as
${\mathbf{G}}_{s}\;=\;-\frac{1}{2}\,(\mathbf{I}_{M+1}-\mathbf{1}_{M+1}\mathbf{a}_{M+1}^{\textrm{T}})\,{\mathbf{D}}_{s}\,(\mathbf{I}_{M+1}-\mathbf{a}_{M+1}\mathbf{1}_{M+1}^{\textrm{T}})\;,$ (12)
where the centering vector is defined as $\mathbf{a}_{M+1}=\frac{1}{M}[\mathbf{1}_{M}^{\textrm{T}},0]^{\textrm{T}}$, again to ensure that the centroid of the relative microphone positions is at the origin. Let us consider the eigenvalue decomposition
${\mathbf{G}}_{s}\;=\;\mathbf{S}\boldsymbol{\Lambda}\mathbf{S}^{\textrm{T}}\;,$ (13)
with $\mathbf{S}$ and $\boldsymbol{\Lambda}$ denoting the matrices containing the eigenvectors and eigenvalues, respectively. Since the rank of the positive semi-definite matrix $\mathbf{G}_{s}$ is at most $P$, this means that at most $P$ eigenvalues are non-zero, i.e., $\lambda_{1}\geq\dots\geq\lambda_{P}\geq 0$ and $\lambda_{P+1}=\dots=\lambda_{M+1}=0$. Using (7) and (13), the relative microphones and source positions matrix $\mathbf{P}_{\textrm{r}s}$ can hence be written as
$\mathbf{P}_{\textrm{r}s}\;=\;\left[\textrm{diag}\left(\sqrt{\lambda_{1}},\dots,\sqrt{\lambda_{P}}\right)\,,\;\mathbf{0}_{{P}\times(M+1-P)}\right]\mathbf{S}^{\textrm{T}}\;.$ (14)
Using (5), the (absolute) source position vector of the $s$-th source $\mathbf{p}_{s}$ can be obtained from $\mathbf{P}_{\textrm{r}s}$ as
$\mathbf{p}_{s}\;=\;\underbrace{\mathbf{R}\mathbf{P}_{\textrm{r}s}}_{\mathbf{P}_{s}}\overline{\mathbf{e}}_{M+1}\;,$ (15)
with $\overline{\mathbf{e}}_{m}$ the $(M+1)$-dimensional selection vector, consisting of zeros except for a one in the $m$-th entry.
III Euclidean Distance Matrix-Based Position Estimation
In Section III-A, we review the EDM-based method in [25] to estimate the 3D position of a single source. Considering multiple candidate TDOA estimates, this method estimates the distance between the source and a reference microphone by minimizing an EDM-based cost function in a single continuous variable, considering all possible combinations of candidate TDOA estimates. In Section III-B, we propose an extension of this method to estimate the 3D positions of multiple sources. The proposed method estimates the optimal set of candidate TDOA estimates and corresponding source distances to the reference microphone by considering the $S$ smallest values of the EDM-based cost function. For each source, the absolute position is then computed from the estimated relative position by solving an orthogonal Procrustes problem.
III-A Single-Source Position Estimation
In this section, we consider a single-source scenario, i.e., $S=1$. The EDM-based method in [25] estimates the source position $\mathbf{p}_{1}$ by first estimating the vector of squared distances $\mathbf{d}_{1}$ between the source and all microphones. To this end, an EDM-based cost function using the eigenvalues of the Gram matrix of the source is defined and minimized. In Section III-A1 we define the cost function assuming the TDOAs to be known. Aiming at improving robustness against noise and reverberation, in Section III-A2 we explain how to incorporate multiple TDOA estimates.
III-A1 EDM-Based Cost Function
By introducing the variable $\alpha$, which represents the distance between the source and the reference microphone, the distance between the source and the $m$-th microphone can be written using (9) as a function of the variable $\alpha$ as
$d_{m1}(\alpha)\;=\;\alpha\;+\;\nu\,\tau_{m}(\mathbf{p}_{1})\;\;\,$, (16)
assuming the TDOAs $\tau_{m}(\mathbf{p}_{1}),\;m=2,\dots,M$, to be known (see Fig. 1). Using (16), we construct the vector of squared distances as $\mathbf{d}_{1}(\alpha)=\left[d^{2}_{11}(\alpha),d^{2}_{21}(\alpha),\dots,d^{2}_{M1}(\alpha)\right]^{\textrm{T}}$ and the EDM in (10) as a function of $\alpha$, i.e.,
${\mathbf{D}}_{1}(\alpha)\;=\;\begin{bmatrix}\begin{array}[]{c|c}\mathbf{D}_{MM}&\mathbf{d}_{1}(\alpha)\\ \hline\cr\mathbf{d}_{1}^{\textrm{T}}(\alpha)&0\end{array}\end{bmatrix}\;$. (17)
The Gram matrix ${\mathbf{G}}_{1}(\alpha)$ corresponding to the EDM ${\mathbf{D}}_{1}(\alpha)$ in (17) can be constructed using (12). Based on the fact that the rank of the Gram matrix $\mathbf{G}_{1}(\alpha)$ attains its minimum value (at most $P$) when $\alpha=d_{11}$, and is larger for other values of $\alpha$, it was proposed in [25] to define a cost function based on all but the $P$ largest eigenvalues $\lambda_{i}(\alpha)$ of $\mathbf{G}_{1}(\alpha)$, i.e.,
$\boxed{J\left(\alpha\right)\;\;\;=\sum_{i=P+1}^{M+1}|\lambda_{i}\left(\alpha\right)|}$ (18)
It should be noted that since it cannot be guaranteed that the eigenvalues of ${\mathbf{G}}_{1}(\alpha)$ are positive for all values of $\alpha$ (e.g., in case of a mismatch between the distance variable and the TDOAs), the absolute values of the eigenvalues are used in (18). If $\alpha=d_{11}$, all but the $P$ largest eigenvalues of ${\mathbf{G}}_{1}(d_{11})$ are equal to zero, such that $J\left(d_{11}\right)=0$. The optimal value $\alpha_{\textrm{opt}}=d_{11}$ can hence be found as
$\alpha_{\textrm{opt}}\;=\;\;\operatorname*{argmin}_{\alpha}\;\;J\left(\alpha\right)\;.$ (19)
It should be noted that no closed-form solution is available, such that an exhaustive search over the (single) continuous variable $\alpha$ needs to be performed.
III-A2 TDOA Estimation and Selection
Since in practice the TDOAs of the source are obviously not available, in [25] a method was proposed to incorporate TDOA estimates into the EDM-based cost function. A commonly used method to estimate TDOAs between microphone pairs is based on the generalized cross correlation with phase transform (GCC-PHAT) function [34]. The continuous-time GCC-PHAT function between the $m$-th and the reference microphone is defined as
$\xi_{m}(\tau)\;=\int_{-\omega_{0}}^{\omega_{0}}\psi_{m1}(\omega)e^{-\jmath\omega\tau}d\omega\;,$ (20)
with radial frequency $\omega$, $\jmath=\sqrt{-1}$, and time lag $\tau$. The normalized phase spectrum $\psi_{m1}(\omega)$ in (20) is given by
$\psi_{m1}(\omega)\;=\;\frac{\operatorname{\mathbb{E}}\{Y_{m}(\omega)Y_{1}^{*}(\omega)\}}{|\operatorname{\mathbb{E}}\{Y_{m}(\omega)Y_{1}^{*}(\omega)\}|}\;,$ (21)
where $Y_{m}(\omega)$ denotes the $m$-th microphone signal in the frequency-domain, $\{\cdot\}^{*}$ denotes the complex-conjugate operator, and $\operatorname{\mathbb{E}}\{\cdot{}\}$ denotes the expectation operator. The TDOA between the $m$-th microphone and the reference microphone is then estimated by determining the main peak of $\xi_{m}(\tau)$, i.e.,
$\hat{\tau}_{m}\;=\;\;\operatorname*{argmax}_{\tau}\;\;\xi_{m}(\tau)\;.$ (22)
Although the PHAT weighting in (21) has been shown to improve robustness against reverberation and noise [35, 36, 37], acoustic reflections can introduce peaks in $\xi_{m}(\tau)$ that are higher than the peak corresponding to the direct source component. Therefore, it was proposed in [25] to consider $C$ candidate TDOA estimates instead of considering only the global maximum of the GCC-PHAT function as in (22). For each microphone $m=2,\dots,M$, the set of $C$ candidate TDOA estimates is denoted as $\hat{\mathcal{T}}_{m}=\{\hat{\tau}_{m}(1),\dots,\hat{\tau}_{m}(C)\}$, corresponding to the $C$ largest local maxima of the GCC-PHAT function $\xi_{m}(\tau)$. To determine which combination of candidate TDOA estimates best fits the source and microphones geometry, we consider all $Q=C^{M-1}$ possible combinations and define the combination vectors
$\mathbf{c}(q)\;=\;\left[\,c_{2}(q),\dots,c_{M}(q)\,\right]\;\;\;,\quad\;q=1,\dots,Q\;,$ (23)
where the index $c_{m}(q)\in\{1,\dots,C\}$ refers to one of the $C$ largest local maxima of $\xi_{m}(\tau)$. For each possible combination of candidate TDOA estimates, the vector of squared distances between the source and the reference microphone is given by
$\mathbf{d}_{1}(\alpha,q)\;=\;\left[d_{11}^{2}(\alpha,q),d_{21}^{2}(\alpha,q),\dots,d_{M1}^{2}(\alpha,q)\right]^{\textrm{T}}\;\;,\quad q=1,\dots,Q\;,$ (24)
where, similarly to (16),
$d_{m1}(\alpha,q)\;=\;\alpha+\nu\hat{\tau}_{m}(c_{m}(q))\;\;,\quad m=1,\dots,M\,,\;q=1,\dots,Q\;.$ (25)
Using (24), for each possible combination of candidate TDOA estimates we can construct the cost function in (18) as
$J(\alpha,q)\;\;\;=\sum_{i=P+1}^{M+1}|\lambda_{i}(\alpha,q)|\;\;,\quad q=1,\dots,Q\;,$ (26)
with $\lambda_{i}(\alpha,q)$ the eigenvalues of the Gram matrix $\mathbf{G}_{1}(\alpha,q)$ associated with the EDM
${\mathbf{D}}_{1}(\alpha,q)\;=\;\begin{bmatrix}\begin{array}[]{c|c}\mathbf{D}_{MM}&\mathbf{d}_{1}(\alpha,q)\\ \hline\cr\mathbf{d}_{1}^{\textrm{T}}(\alpha,q)&0\end{array}\end{bmatrix}\;\;,\quad q=1,\dots,Q\;.$ (27)
The optimal combination of candidate TDOA estimates is then determined by first determining the optimal distance variable for each combination, i.e.,
$\hat{\alpha}(q)\;=\;\;\operatorname*{argmin}_{\alpha}\;\;J(\alpha,q)\;\;,\quad q=1,\dots,Q\;,$ (28)
and then determining the combination which results in the smallest value of the cost function at these solutions, i.e.,
$\hat{q}_{1}\;=\;\operatorname*{argmin}_{q=1,\dots,Q}\;J(\hat{\alpha}(q),q)\;,$ (29)
with corresponding Gram matrix $\hat{\mathbf{G}}_{1}={\mathbf{G}}_{1}(\hat{\alpha}(\hat{q}_{1}),\hat{q}_{1})$. It should be noted that due to possible TDOA estimation errors, $J(\hat{\alpha}(\hat{q}_{1}),\hat{q}_{1})$ is not guaranteed to be 0 unlike $J({\alpha}_{\textrm{opt}})$ from (19).
As explained in Section II, from the eigenvalue decomposition of the estimated Gram matrix $\hat{\mathbf{G}}_{1}=\hat{\mathbf{S}}\hat{\boldsymbol{\Lambda}}\hat{\mathbf{S}}^{\textrm{T}}$, the relative microphones and source positions matrix can be estimated as in (14) using the $P$ largest eigenvalues, i.e.,
$\hat{\mathbf{P}}_{\textrm{r}1}\;=\;\left[\textrm{diag}\left(\sqrt{\hat{\lambda}_{1}},\dots,\sqrt{\hat{\lambda}_{P}}\right)\,,\;\mathbf{0}_{{P}\times(M+1-P)}\right]\hat{\mathbf{S}}^{\textrm{T}}\;.$ (30)
It should be noted that due to estimation errors it cannot be guaranteed that the estimated relative microphone positions matrix $\hat{\mathbf{M}}_{\textrm{r}1}$ (first $M$ columns of $\hat{\mathbf{P}}_{\textrm{r}1}$) can be perfectly mapped to the known absolute microphone position matrix $\mathbf{M}$ by rotation and/or reflection as in (4). As proposed in [25], the mapping between the estimated relative microphone positions $\hat{\mathbf{M}}_{\textrm{r}1}$ and the absolute microphone positions $\mathbf{M}$ can be computed by solving an orthogonal Procrustes problem [38, 33]. First, the singular value decomposition (SVD) of $\hat{\mathbf{M}}_{\textrm{r}1}\mathbf{M}^{\textrm{T}}$ is computed as
$\hat{\mathbf{M}}_{\textrm{r}1}\mathbf{M}^{\textrm{T}}\;=\;\hat{\mathbf{U}}_{1}\hat{\mathbf{Q}}_{1}\hat{\mathbf{V}}_{1}^{\textrm{T}}\;,$ (31)
where $\hat{\mathbf{Q}}_{1}$ contains the singular values and $\hat{\mathbf{U}}_{1}$ and $\hat{\mathbf{V}}_{1}$ contain the left and right singular vectors, respectively. The orthogonal mapping matrix $\hat{\mathbf{R}}_{1}$ in (5), is then computed using the orthogonal Procrustes problem solution as
$\hat{\mathbf{R}}_{1}\;=\;\hat{\mathbf{V}}_{1}\hat{\mathbf{U}}_{1}^{\textrm{T}}\;,$ (32)
and the absolute source position is then computed using (15) as
$\hat{\mathbf{p}}_{1}\;=\;\hat{\mathbf{R}}_{1}\hat{\mathbf{P}}_{\textrm{r}1}\overline{\mathbf{e}}_{M+1}\;.$ (33)
III-B Multi-Source Position Estimation
In this section, we propose an extension of the EDM-based method presented in the previous section, to estimate the position of multiple sources. For each source, we now introduce the variable $\alpha_{s}$, which represents the distance between the $s$-th source and the reference microphone. Similarly as in (16), the distance between the $s$-th source and the $m$-th microphone can be written as a function of the variable $\alpha_{s}$ as $d_{ms}(\alpha_{s})=\alpha_{s}+\nu\tau_{m}(\mathbf{p}_{s})\,,\,\;\;\;m=2,\dots,M\,,\,\;\;\;s=1,\dots,S$. Using $\mathbf{d}_{s}(\alpha_{s})\;=\;\left[d_{1s}^{2}(\alpha_{s}),d_{2s}^{2}(\alpha_{s}),\dots,d_{Ms}^{2}(\alpha_{s})\right]^{\textrm{T}}$, we can construct the vector of squared distances for the $s$-th source.
It is indeed possible to construct an $(M+S)\times(M+S)$-dimensional EDM for all sources, similarly to (17), i.e.,
[…]
where it should be realized that this EDM depends on $S$ continuous distance variables and furthermore contains the $S\times S$-dimensional EDM $\mathbf{D}_{SS}$ with unknown inter-source distances. However, since jointly optimizing $\alpha_{1},\dots,\alpha_{S}$ and the unknown entries of $\mathbf{D}_{SS}$ is not straightforward, we propose a simpler procedure, assuming that each source corresponds to a unique combination of TDOAs.
Similarly as for single-source position estimation, we consider $C$ TDOA estimates for the microphones $m=2,\dots,M$ with $C\geq S$. For each combination of candidate TDOA estimates, the optimal value $\hat{\alpha}(q)\,,\;q=1,\dots,Q,$ is determined using (28). Instead of only considering the overall smallest value of the cost function at these solutions, as in the single-source case, we now consider the $S$ smallest values $J(\hat{\alpha}(\hat{q}_{1}),\hat{q}_{1}),\dots,J(\hat{\alpha}(\hat{q}_{2}),\hat{q}_{S})$ and their corresponding Gram matrices $\hat{\mathbf{G}}_{1},\dots,\hat{\mathbf{G}}_{S}$. For each source, the relative microphones and source positions matrix $\hat{\mathbf{P}}_{\textrm{r}s}$ is estimated based on the eigenvalue decomposition of the corresponding Gram matrix $\hat{\mathbf{G}}_{s}$, similarly to (30). Then, similarly to (33), the position of the $s$-th source is then estimated as
$\hat{\mathbf{p}}_{\textrm{r}s}\;=\;\hat{\mathbf{R}}_{s}\hat{\mathbf{P}}_{\textrm{r}s}\overline{\mathbf{e}}_{M+1}\;,$ (35)
with $\hat{\mathbf{R}}_{s}$ the orthogonal mapping matrix for the $s$-th source, computed using the left and right singular vectors $\hat{\mathbf{M}}_{\textrm{r}s}\mathbf{M}^{\textrm{T}}$.
Since for reverberant and noisy scenarios with multiple sources the peaks of the GCC-PHAT function may sometimes contain spurious peaks that don’t correspond to the true TDOAs, it is often beneficial to set $C>S$.
For an exemplary scenario with $M=6$ spatially distributed microphones, $S=2$ sources (at distances $d_{11}=0.95$ m and $d_{12}=2.46$ m from the reference microphone) and $C=2$ candidate TDOA estimates, Fig. 2(a) illustrates the cost function $J(\alpha,q)$ in (26) for all $Q=32$ combinations of candidate TDOA estimates, while Fig. 2(b) shows the sorted minimum cost function values $J(\hat{\alpha}(q),q)$. For this scenario, it can be observed that only 2 out of 32 cost functions exhibit clear minima. Moreover, it can be observed that the estimated distance variables $\hat{\alpha}_{1}$ and $\hat{\alpha}_{2}$ for these combinations of TDOA estimates correspond very closely to the distances $d_{11}$ and $d_{12}$.
Figure 2: (a) Cost functions $J(\alpha,q)$ for all combinations of TDOA estimates $q$, (b) Corresponding minimum cost function values.
IV Euclidean Distance Matrix-Based DOA Estimation
Whereas in Section III we considered spatially distributed microphones and proposed an EDM-based method to estimate the 3D positions of multiple sources, in this section we will consider compact microphone arrays and assume that the sources are in the far field of the microphone array. In Section IV-A we propose an EDM-based method to estimate the DOA of a single source. By defining a relative coordinate system in which one of the basis vectors is the DOA vector of the source, the rank of the Gram matrix associated with the EDM containing the microphone distances can be reduced by subtracting a rank-1 matrix based on the TDOAs. Using the eigenvalues of this rank-reduced Gram matrix, we define an EDM-based cost function to estimate the DOA, which does not require continuous variable optimization. In Section IV-B we extend this method to estimate the DOAs of multiple sources. Similarly as for multi-source position estimation, we determine the optimal set of candidate TDOA estimates based on the EDM-based cost function, solving the association of candidate TDOA estimates to sources.
IV-A Single-Source DOA Estimation
We consider a compact microphone array and a source in the far field of the microphone array, assuming that the acoustic waves arriving at the microphones from the source can be approximated as planar waves (see Fig. 3 for an exemplary two-dimensional configuration). This assumption is valid when the distances between the source and the microphones are much larger than the inter-microphone distances. In Section IV-A1 we define a rank-reduced Gram matrix, assuming the TDOAs to be known. In Section IV-A2 we explain how to incorporate multiple TDOA estimates.
IV-A1 Rank-Reduced Gram Matrix
In the absolute 3D coordinate system (with canonical basis vectors $\mathbf{e}_{1}$, $\mathbf{e}_{2}$, and $\mathbf{e}_{3}$), the unit-norm DOA vector, pointing in the direction of the source (see Fig. 3),
[…]
Figure 3: Exemplary two-dimensional configuration, consisting of a compact microphone array with $M=4$ microphones and a far-field source at azimuth angle $\theta_{1}$. In the relative coordinate system, defined by the DOA vector $\mathbf{v}_{1}$, the relative distances between the $m$-th microphone and the centroid of the microphone array, in the direction of the source, are denoted by $\delta_{m}\,,\,\;\;m=1,\dots,M$.
is defined as
$\mathbf{v}_{1}\;=\;[\textrm{cos}(\theta_{1})\textrm{cos}(\phi_{1})\;,\;\;\;\textrm{sin}(\theta_{1})\textrm{cos}(\phi_{1})\;,\;\;\;\textrm{sin}(\phi_{1})]^{\textrm{T}}\;,$ (36)
with $\theta_{1}$ and $\phi_{1}$ the azimuth and elevation angle, respectively. These angles can be computed from the DOA vector as
$\theta_{1}\;$ $=\;\textrm{tan}^{-1}\left(\frac{\mathbf{e}_{2}^{\textrm{T}}\mathbf{v}_{1}}{\mathbf{e}_{1}^{\textrm{T}}\mathbf{v}_{1}}\right)\;,$ (37)
$\phi_{1}\;$ $=\;\sin^{-1}\left(\mathbf{e}_{3}^{\textrm{T}}\mathbf{v}_{1}\right)\;.$ (38)
We now define a relative coordinate system with orthonormal basis vectors $\mathbf{e}_{\textrm{x}}$, $\mathbf{e}_{\textrm{y}}$, and $\mathbf{e}_{\textrm{z}}$, which are not necessarily canonical, where we set one of the basis vectors equal to the DOA vector, e.g., $\mathbf{e}_{\textrm{x}}=\mathbf{v}_{1}$. The basis vectors of the relative coordinate system can be mapped to the basis vectors of the absolute coordinate system through a rotation with the angles $\theta_{1}$ and $\phi_{1}$, i.e.,
$\mathbf{e}_{\textrm{x}}\;=\;\mathbf{v}_{1}\;=\;\mathbf{R}_{1}\mathbf{e}_{1}\quad,\;\quad\mathbf{e}_{\textrm{y}}\;=\;\mathbf{R}_{1}\mathbf{e}_{2}\quad,\;\quad\mathbf{e}_{\textrm{z}}\;=\;\mathbf{R}_{1}\mathbf{e}_{3}\;,$ (39)
with orthogonal mapping matrix
$\mathbf{R}_{1}=\begin{bmatrix}\cos(\theta_{1})\cos(\phi_{1})&-\sin(\theta_{1})&-\cos(\theta_{1})\sin(\phi_{1})\\ \sin(\theta_{1})\cos(\phi_{1})&\cos(\theta_{1})&-\sin(\theta_{1})\sin(\phi_{1})\\ \sin(\phi_{1})&0&\cos(\phi_{1})\end{bmatrix}\;$. (40)
Using this mapping matrix, the microphone positions in the relative coordinate system (39) can be written in the absolute coordinate system, i.e., the relative microphone positions matrix $\mathbf{M}_{\textrm{r}}=\mathbf{R}_{1}^{\textrm{T}}\mathbf{M}$. The $M$-dimensional relative coordinate vectors $\mathbf{x}_{\textrm{r}}$, $\mathbf{y}_{\textrm{r}}$, and $\mathbf{z}_{\textrm{r}}$ are defined as
$\mathbf{M}_{\textrm{r}}\;=\;\begin{bmatrix}\mathbf{x}_{\textrm{r}}^{\textrm{T}}\\ \mathbf{y}_{\textrm{r}}^{\textrm{T}}\\ \mathbf{z}_{\textrm{r}}^{\textrm{T}}\end{bmatrix}\;$. (41)
It should be realized that the coordinate vector $\mathbf{x}_{\textrm{r}}$ contains the relative distances between the microphones and the centroid of the microphone array in the direction of the source (see Fig. 3). Since the relative distance $\delta_{m}$ between the $m$-th microphone and the centroid of the microphone array is directly related to the TDOA $\tilde{\tau}_{m}=\mathbf{m}_{m}^{\textrm{T}}\mathbf{v}_{1}/\nu$ between the $m$-th microphone and the centroid of the microphone array as $\delta_{m}=-\nu\tilde{\tau}_{m}$, the coordinate vector $\mathbf{x}_{\textrm{r}}$ can be written as
$\mathbf{x}_{\textrm{r}}\;=\;-\nu\tilde{\boldsymbol{\tau}}\;$, (42)
where $\tilde{\boldsymbol{\tau}}=\left[\tilde{\tau}_{1},\dots,\tilde{\tau}_{M}\right]^{\textrm{T}}$ denotes the vector of (centered) TDOAs, for which $\tilde{\boldsymbol{\tau}}^{\textrm{T}}\mathbf{1}_{M}=0$.
Using (41), it can easily be seen that the $M\times M$-dimensional Gram matrix of the microphones ${\mathbf{G}}_{MM}$ in (6) can be written as the sum of three rank-1 matrices, i.e.,
$\mathbf{G}_{MM}\;=\;\mathbf{x}_{\textrm{r}}\mathbf{x}_{\textrm{r}}^{\textrm{T}}+\mathbf{y}_{\textrm{r}}\mathbf{y}_{\textrm{r}}^{\textrm{T}}+\mathbf{z}_{\textrm{r}}\mathbf{z}_{\textrm{r}}^{\textrm{T}}\;,$ (43)
whose rank is at most 3 (i.e., $P$). Assuming the TDOA vector $\tilde{\boldsymbol{\tau}}$ to be known, we now define the Gram matrix
$\boxed{\mathbf{G}_{MM}^{\;-}\;=\;\mathbf{G}_{MM}\;-\;\nu^{2}\tilde{\boldsymbol{\tau}}\tilde{\boldsymbol{\tau}}^{\textrm{T}}}$ (44)
Using (42) and (43), it can be easily seen that $\mathbf{G}_{MM}^{\;-}=\mathbf{y}_{\textrm{r}}\mathbf{y}_{\textrm{r}}^{\textrm{T}}+\mathbf{z}_{\textrm{r}}\mathbf{z}_{\textrm{r}}^{\textrm{T}}$, whose rank is at most 2 (i.e., P1P-1). Considering the eigenvalue decomposition
$\mathbf{G}_{MM}^{\;-}=\mathbf{W}\boldsymbol{\Sigma}\mathbf{W}^{\textrm{T}}\;,$ (45)
with the matrix $\mathbf{W}$ containing the eigenvectors and the diagonal matrix $\boldsymbol{\Sigma}$ containing the eigenvalues, this means that at most $2$ eigenvalues are positive, $\sigma_{1}\geq\sigma_{2}\geq 0$, while the other eigenvalues are equal to zero, $\sigma_{3}=\dots=\sigma_{M}=0$. By defining
$\mathbf{M}_{\textrm{r}}^{\;-}\;=\;\begin{bmatrix}\mathbf{y}_{\textrm{r}}^{\textrm{T}}\\ \mathbf{z}_{\textrm{r}}^{\textrm{T}}\end{bmatrix}\;$, (46)
the rank-reduced Gram matrix can be written as $\mathbf{G}_{MM}^{\;-}=(\mathbf{M}_{\textrm{r}}^{\;-})^{\textrm{T}}\mathbf{M}_{\textrm{r}}^{\;-}$. Based on (14), the matrix $\mathbf{M}_{\textrm{r}}^{\;-}$ can be computed from (45) up to an arbitrary orthogonal transformation $\mathbf{R}_{\textrm{ar}}^{\;-}$, i.e.,
$\mathbf{M}_{\textrm{ar}}^{-}\;=\;\left[\operatorname{\rm{diag}}\left(\sqrt{\sigma_{1}},\sqrt{\sigma_{2}}\right)\,,\;\mathbf{0}_{2\times(M-2)}\right]\mathbf{W}^{\textrm{T}}\;,$ (47)
with $\mathbf{M}_{\textrm{r}}^{-}=\mathbf{R}_{\textrm{ar}}^{-}\mathbf{M}_{\textrm{ar}}^{-}$. Using (42) and (46), the relative microphone positions can be written as
[…]
Therefore, using (4), the absolute microphone positions can be written as
$\mathbf{M}\;=\;\mathbf{R}_{1}\mathbf{M}_{\textrm{r}}\;=\;\underbrace{\mathbf{R}_{1}\mathbf{R}_{\textrm{ar}}}_{\mathbf{R}^{\prime}}\mathbf{M}_{\textrm{ar}}\;,$ (49)
with $\mathbf{R}^{\prime}$ an orthogonal mapping matrix. Realizing that the canonical basis vector $\mathbf{e}_{1}$ of the absolute coordinate system is equal to the first column of the matrix $\mathbf{R}_{\textrm{ar}}$ defined in (48), i.e., $\mathbf{e}_{1}=\mathbf{R}_{\textrm{ar}}\mathbf{e}_{1}$, the DOA vector $\mathbf{v}_{1}$ in (39) can be written as
$\boxed{\mathbf{v}_{1}\;=\;\mathbf{R}_{1}\mathbf{R}_{\textrm{ar}}\mathbf{e}_{1}\;=\;\mathbf{R}^{\prime}\mathbf{e}_{1}\;}$ (50)
This means that the DOA vector is equal to the first column of the matrix $\mathbf{R}^{\prime}$ mapping the matrix $\mathbf{M}_{\textrm{ar}}$ to the absolute microphone positions matrix $\mathbf{M}$, which can be determined by solving an orthogonal Procrustes problem (without explicitly needing to compute the mapping matrices $\mathbf{R}_{1}$ and $\mathbf{R}_{\textrm{ar}}^{-}$).
IV-A2 TDOA Estimation and Selection
Computing the matrices $\mathbf{G}_{MM}^{\;-}$ in (44) and $\mathbf{M}_{\textrm{ar}}$ in (48) requires centered TDOAs $\tilde{\boldsymbol{\tau}}_{m}$, which are not available in practice. Similarly as in Section III-A2, we consider $C$ candidate TDOA estimates $\hat{\tilde{\boldsymbol{\tau}}}_{m}\;,$$m=2,\dots,M$, corresponding to the $C$ largest local maxima of the GCC-PHAT function, and consider the $Q=C^{M-1}$ combination vectors $\mathbf{c}(q)$ in (23). For each combination $q$ of candidate TDOA estimates, the centered TDOA between the $m$-th microphone and the centroid of the microphone array can be estimated as
$\hat{\widetilde{\tau}}_{m}(q)\;=\;\hat{\tau}_{m}(c_{m}(q))-\frac{1}{M}\sum_{j=1}^{M}\hat{\tau}_{j}(c_{j}(q))\;.$ (51)
The selection of candidate TDOA estimates will be explained in the next section for multiple sources (which can obviously also be used for a single source).
IV-B Multi-Source DOA Estimation
In this section, we present a method to estimate the DOAs of multiple sources based on the eigenvalues of the Gram matrix defined in (44). Contrary to the EDM-based position estimation method in Section III, which requires an exhaustive search over the continuous variable $\alpha$, the proposed EDM-based DOA estimation method does not require continuous variable optimization.
Using the vector of centered TDOA estimates $\hat{\tilde{\boldsymbol{\tau}}}(q)=[\hat{\widetilde{\tau}}_{1}(q),\dots,\hat{\widetilde{\tau}}_{M}(q)]^{\textrm{T}}$ based on (51), the Gram matrix in (44) is defined for each possible combination of candidate TDOA estimates as
$\hat{\mathbf{G}}_{MM}^{\;-}(q)\;=\;{\mathbf{G}}_{MM}-\nu^{2}\hat{\tilde{\boldsymbol{\tau}}}(q)\hat{\tilde{\boldsymbol{\tau}}}^{\textrm{T}}(q)\;\;\;,\quad q=1,\dots,Q\;.$ (52)
Similarly to (26), we define a cost function based on all but the P1P-1 largest eigenvalues $\hat{\sigma}_{i}(q)$ of $\hat{\mathbf{G}}_{MM}^{\;-}(q)$, i.e.,
$\boxed{I\left(q\right)\;\;=\sum_{i=P}^{M}|\hat{\sigma}_{i}(q)|\;\;\;\;,\quad q=1,\dots,Q}$ (53)
For a single source ($S=1$), the optimal combination of candidate TDOA estimates is determined as the one minimizing the cost function, i.e.,
$\hat{q}_{1}\;=\;\;\operatorname*{argmin}_{q=1,\dots,Q}\;\;I\left(q\right)\;.$ (54)
It should be noted that due to possible TDOA estimation errors the rank of $\hat{\mathbf{G}}_{MM}^{\;-}(\hat{q}_{1})$ is not guaranteed to be equal to P1P-1. For multiple sources, we consider the $S$ smallest cost function values $I(\hat{q}_{1}),\dots,I(\hat{q}_{S})$ and their corresponding Gram matrices $\hat{\mathbf{G}}_{MM}^{\;-}(\hat{q}_{s})$. Using the eigenvalue decomposition $\hat{\mathbf{G}}_{MM}^{\;-}(\hat{q}_{s})=\hat{\mathbf{W}}(\hat{q}_{s})\hat{\boldsymbol{\Sigma}}(\hat{q}_{s})\hat{\mathbf{W}}^{\textrm{T}}(\hat{q}_{s})$, the matrices $\mathbf{M}_{\textrm{ar}}^{-}$ in (47) and $\mathbf{M}_{\textrm{ar}}$ in (48) can be estimated for each source as
$\begin{split}\hat{\mathbf{M}}_{\textrm{ar}}^{-}(\hat{q}_{s})\;&=\\ &\hskip-38.41139pt\bigl[\textrm{diag}\left(\sqrt{\hat{\sigma}_{1}(\hat{q}_{s})},\dots,\sqrt{\hat{\sigma}_{P-1}(\hat{q}_{s})}\right)\,,\;\mathbf{0}_{{(P-1)}\times(M-P+1)}\bigr]\,\hat{\mathbf{W}}^{\textrm{T}}(\hat{q}_{s}),\end{split}$ (55)
[…]
The mapping matrix $\hat{\mathbf{R}}^{\prime}(\hat{q}_{s})$ between the matrix $\hat{\mathbf{M}}_{\textrm{ar}}(\hat{q}_{s})$ and the absolute microphone positions matrix $\mathbf{M}$ can be computed by solving an orthogonal Procrustes problem, similarly to (31) and (32). From these mapping matrices, the DOA vectors can then be estimated using (50) as $\hat{\mathbf{v}}_{s}=\hat{\mathbf{R}}^{\prime}(\hat{q}_{s})\;\mathbf{e}_{1}$.
For an exemplary scenario with $M=6$ closely spaced microphones, $S=2$ sources (at distances $d_{\textrm{c}1}=1$ m and $d_{\textrm{c}2}=2$ m from the centroid of the microphone array) and $C=2$ candidate TDOA estimates, Fig. 4 shows the sorted cost function values I(q)I({q}) for all $Q=32$ combinations of TDOA estimates. For this scenario, it can be observed that there is a clear difference between the two smallest values and the other values.
Figure 4: Cost function values $I(q)$ for all combinations of candidate TDOA estimates $q$.
V Baseline Localization Methods
In this section, we briefly review SRP-based methods to localize multiple sources, both for position estimation (Section V-A) and DOA estimation (Section V-B). These methods will be used as baseline methods for the experimental evaluation in the next section. Contrary to the EDM-based methods proposed in Sections III and IV, SRP-based methods require joint optimization of a functional over multiple continuous variables.
V-A SRP-Based Multi-Source Position Estimation
Similarly to the GCC-PHAT function in (20), the SRP-PHAT functional for the 3-dimensional position estimation [22] is defined as
$\Psi(\mathbf{p})\;\;=\sum_{i>j}\int_{-\omega_{0}}^{\omega_{0}}\psi_{ij}(\omega)e^{-\jmath\omega\tau_{ij}(\mathbf{p})}\,d\omega\;,$ (57)
where $\tau_{ij}(\mathbf{p})$ denotes the TDOA (8) corresponding to position $\mathbf{p}=[p_{x},p_{y},p_{z}]^{\textrm{T}}$, and the summation considers all microphone pairs, where $i>j$. The estimated source positions $\hat{\mathbf{p}}_{1},\dots,\hat{\mathbf{p}}_{S}$ are determined as the vectors that correspond to the $S$ largest local maxima of the SRP-PHAT functional. This requires a joint optimization of three continuous variables for all feasible positions within the room boundaties ($P_{x}^{\textrm{min}}\leq p_{x}\leq P_{x}^{\textrm{max}}$, $P_{y}^{\textrm{min}}\leq p_{y}\leq P_{y}^{\textrm{max}}$, and $P_{z}^{\textrm{min}}\leq p_{z}\leq P_{z}^{\textrm{max}}$). Practical considerations regarding the discretization of the continuous position variables and the determination of local maxima are discussed in Section VI.
V-B SRP-Based Multi-Source DOA Estimation
Similarly to (57), the SRP-PHAT functional for DOA estimation is defined as
$\Psi(\mathbf{v})\;\;=\sum_{i>j}\int_{-\omega_{0}}^{\omega_{0}}\psi_{ij}(\omega)e^{-\jmath\omega\tau_{ij}(\mathbf{v})}\,d\omega\,$, (58)
where $\tau_{ij}(\mathbf{v})=(\mathbf{m}_{i}-\mathbf{m}_{j})^{\text{T}}\mathbf{v}/\nu$ denotes the TDOA corresponding to the DOA vector $\mathbf{v}$, which depends on the azimuth and elevation angles. The estimated azimuth and elevation angles $\hat{\theta}_{1},\dots,\hat{\theta}_{S}$ and $\hat{\phi}_{1},\dots,\hat{\phi}_{S}$ are determined from the DOA vectors that correspond to the $S$ largest local maxima of the SRP-PHAT functional. This requires a joint optimization of two continuous variables for all feasible azimuth angles $\pi<\theta\leq\pi$ and elevation angles $\pi/2<\phi\leq\pi/2$. Practical considerations regarding the discretization of the azimuth and elevation angles and the determination of local maxima are discussed in Section VI.
VI Experimental Evaluation
In this section, we compare the performance of the proposed EDM-based methods with the baseline SRP-based methods for multi-source position and DOA estimation in noisy and reverberant environments. We conducted two sets of experiments, where in experiment 1 we evaluated 3D position estimation using spatially distributed microphones and in experiment 2 we evaluated 3D DOA estimation using compact microphone arrays. Section VI-A describes the acoustic scenarios for both experiments. The implementation details of the EDM-based and SRP-based methods are presented in Sections VI-B and VI-C, respectively. The results of experiment 1 (position estimation) are discussed in Section VI-D, while the results of experiment 2 (DOA estimation) are discussed in Section VI-E.
VI-A Acoustic Scenarios
For both experiments, we considered a rectangular room with dimensions 6 m $\times$ 6 m $\times$ 2.4 m, $M=6$ microphones and $S=2$ static speech sources. For each experiment, 100 different acoustic scenarios were simulated, where the room impulse responses (RIRs) between the sources and the microphones were generated using the image source method [39, 40], assuming equal reflection coefficients for all walls.
In experiment 1, the spatially distributed microphones were positioned randomly for each scenario within a cube with sides of length 2 m, with a minimum distance of 10 cm between the microphones. In experiment 2, the microphones were positioned randomly for each scenario within a smaller cube with sides of length 10 cm, with a minimum distance of 4 cm between microphones. In both experiments, the distance between the second source and the centroid of the microphone array was fixed at $d_{\textrm{c}2}=2$ m. To evaluate the influence of source distance on localization accuracy, different distances $d_{\textrm{c}1}$ between the first source and the centroid of the microphone array were considered. In experiment 1, we considered $d_{\textrm{c}1}\in\{0,1,2,3,4\}$ m, where $d_{\textrm{c}1}=0$ m and $d_{\textrm{c}1}=1$ m correspond to the first source being inside of the cube, while $d_{\textrm{c}1}=3$ m and $d_{\textrm{c}1}=4$ m correspond to the first source being outside of the cube. In experiment 2, we considered $d_{\textrm{c}1}\in\{0.5,1,2,3,4\}$ m, corresponding to both sources being outside of the cube (in the far field of the microphone array). For all scenarios, both sources were spaced at least 1 m apart and maintained a minimal angular separation of $20^{\circ}$ relative to the array centroid.
For each scenario, a 5-second speech signal, randomly selected from the M-AILABS dataset [41], with equal probability of being a male or female speaker, was used for each source and convolved with the simulated RIRs. The sampling frequency was equal to $f_{s}=16$ kHz. Spherically isotropic multi-talker babble noise generated using [42] was added to the reverberant speech mixtures at the microphones, with a reverberant signal-to-noise ratio (SNR) of 20 dB (averaged across the microphones). For each scenario, the reflection coefficients were set such that the direct-to-reverberant ratio (DRR) of the second source (at a fixed distance $d_{\textrm{c}2}=2$ m) was equal to 5 dB (averaged across the microphones). The resulting reverberation times were equal to $T_{60}\approx 186\pm 16$ ms for experiment 1 and $T_{60}\approx 193\pm 20$ ms for experiment 2. Obviously, the power ratio of the reverberant speech signals between the first and the source (averaged across microphones and scenarios) decreased for increasing source distance $d_{\textrm{c}1}$, being approximately equal to 0 dB when $d_{\textrm{c}1}=d_{\textrm{c}2}=2$ m.
VI-B Implementation of EDM-Based Source Localization
As explained in Sections III and IV, the proposed EDM-based localization methods require candidate TDOA estimates, which are obtained from the GCC-PHAT function. In practice, the time-domain microphone signals are first transformed to the short-time Fourier transform (STFT) domain, with a frame length of $K=512$ samples (corresponding to 32 ms), $50\%$ overlap between frames, and using a square-root-Hann analysis window. Similarly to (21), the instantaneous normalized phase spectrum between the $i$-th and $j$-th microphones is computed as
$\psi_{ij}[k,l]\;=\;\frac{Y_{i}[k,l]Y_{j}^{*}[k,l]}{|Y_{i}[k,l]Y_{j}^{*}[k,l]|}\;,$ (59)
where $Y_{i}[k,l]$ denotes the STFT coefficient of the $i$-th microphone signal at frequency bin $k\in\{0,\dots,K-1\}$ and time frame $l\in\{1,\dots,L\}$, where $L$ denotes the number of frames. Similarly to (20), the discrete-time GCC-PHAT function between the $m$-th microphone and the reference microphone is computed as
$\xi_{m}[n,l]\;\;=\sum_{k=0}^{K-1}\psi_{m1}[k,l]e^{-\jmath 2\pi nk/K}\;\;\;,\quad\;m=2,\dots,M\;,$ (60)
where $n$ denotes the discrete-time index, with $\tau=n/f_{s}$. To achieve a more precise TDOA estimate, the discrete-time GCC-PHAT function $\xi_{m}[n,l]$ is interpolated by a factor $R=20$ using resampling. Only plausible time-lags $n_{m}$ between the $m$-th microphone and the reference microphone are considered, determined by the inter-microphone distance $D_{m1}$, i.e., $|n_{m}|<Rf_{s}D_{m1}/\nu$. To emphasize strong peaks, the function is weighted as $\tilde{\xi}_{m}[n_{m},l]=\exp{(\gamma\xi_{m}[n_{m},l])}$, with $\gamma=30$ for position estimation (experiment 1) and $\gamma=50$ for DOA estimation (experiment 2). Since the sources are assumed to be static, the GCC-PHAT function is averaged over all frames, i.e.,
$\tilde{\xi}_{m}[n_{m}]\;=\;\frac{1}{L}\sum_{l=1}^{L}\tilde{\xi}_{m}[n_{m},l]\;.$ (61)
The $C$ discrete candidate TDOA estimates $\hat{n}_{m}(1),\dots,\hat{n}_{m}(C)$ between the $m$-th microphone and the reference microphone are determined from (61) by using a peak-finding algorithm, which picks the $C$ highest local maxima. Each estimated TDOA is then fine-tuned using quadratic interpolation [43]. The continuous candidate TDOA estimates are then computed as $\hat{\tau}_{m}(c_{m}(q))=\hat{n}_{m}(c_{m}(q))/(Rf_{s})$.
For EDM-based position estimation, the optimal distance variable in (28) was first determined using a one-dimensional grid search, with a resolution of 1 cm and a maximum distance of 6 m (i.e., 601 grid points) and then fine-tuned with a quadratic interpolation. $C=3$ candidate TDOA estimates are considered per microphone pair, resulting in $Q=C^{M-1}=243$ total combinations. After computing the EDM-based cost functions $J(\alpha,q)$ in (26) and determining the optimal distance variable $\hat{\alpha}(q)$ using a one-dimensional grid search for all $Q$ possible combinations of candidate TDOA estimates, the respective cost function minima were sorted. The combination $\hat{q}_{1}$ yielding the smallest cost function minimum $J(\hat{\alpha}(\hat{q}_{1}),\hat{q}_{1})$ was used to estimate the source position $\hat{\mathbf{p}}_{1}$. To estimate the second source position $\hat{\mathbf{p}}_{2}$, only combinations where at least four candidate TDOA estimates differ were considered, as the likelihood that more than three TDOAs are shared between sources was assumed to be negligible.
For EDM-based DOA estimation, only $C=2$ candidate TDOA estimates were considered, resulting in $Q=32$ total combinations. Unlike with spatially distributed microphones, preliminary experiments with compact microphone arrays indicated that choosing $C>2$ did not notably improve the accuracy of the localization accuracy, since the number of spurious peaks in the GCC-PHAT function is typically quite low when considering plausible time-lags for compact microphone arrays. After computing the EDM-based cost function values $I(q)$ in (53) for all $Q$ possible combinations of candidate TDOA estimates, the respective cost function values were sorted. The combination $\hat{q}_{1}$ yielding the smallest value $I(\hat{q}_{1})$ was used to estimate the source DOA $\hat{\mathbf{v}}_{1}$. To estimate the second source DOA $\hat{\mathbf{v}}_{2}$, only combinations where at least four candidate TDOA estimates differ were considered.
It is known that the choice of reference microphone affects the TDOA estimation accuracy for single source scenarios [44]. Since preliminary experiments using multiple sources also demonstrated that the localization accuracy of the EDM-based methods may be negatively influenced by a poor choice of reference microphone, the reference microphone was chosen depending on the array geometry. For spatially distributed microphone arrays (experiment 1), the reference microphone was chosen as the microphone closest to the array centroid. This reduces the likelihood that the reference microphone is located at a large distance from one or both sources, which would result in a poor SNR. For compact microphone arrays (experiment 2), the reference microphone was chosen as the microphone which was on average farthest from the other microphones, aiming at mitigating the influence of reverberation and noise on the reliability of the GCC-PHAT function [45].
VI-C Implementation of SRP-Based Source Localization
For the SRP-based localization methods (Section V), the exhaustive search over continuous variables is implemented as a discretized grid search. Using the instantaneous normalized phase spectrum in (59), the SRP-PHAT functionals for position and DOA estimation in (57) and (58), are computed as
$\Psi[l]({\mathbf{p}})\;\;\;\;\,={\sum_{i>j}\sum_{k=0}^{K-1}\psi_{ij}[k,l]e^{-\jmath 2\pi f_{s}\tau_{ij}({\mathbf{p}})k/K}}\;\;,$ (62)
$\Psi[l]({\mathbf{v}})\;\;\;\;\,={\sum_{i>j}\sum_{k=0}^{K-1}\psi_{ij}[k,l]e^{-\jmath 2\pi f_{s}\tau_{ij}({\mathbf{v}})k/K}}\;.$ (63)
Similarly as for the GCC-PHAT function in (61), the SRP-PHAT functionals are averaged over all frames, i.e.,
$\Psi({\mathbf{p}})\;=\;\frac{1}{L}\sum_{l=1}^{L}\Psi[l]({\mathbf{p}})\;\;,$ (64)
$\Psi({\mathbf{v}})\;=\;\frac{1}{L}\sum_{l=1}^{L}\Psi[l]({\mathbf{v}})\;\;.$ (65)
Exhaustively searching for the local maxima of these functionals can be computationally demanding if the grid resolution is high (especially for three-dimensional position estimation). Therefore, we first compute $\beta\geq S$ candidate positions/DOAs on a coarse grid (similarly to [46]) that spans the entire range of feasible positions/DOAs. We then refine these coarse estimates by re-evaluating the functional on a finer grid in the vicinity of the coarse estimates. Finally, the $S$ positions $\hat{\mathbf{p}}_{1}^{\textrm{SRP}},\dots,\hat{\mathbf{p}}_{S}^{\textrm{SRP}}$ or DOAs $\hat{\mathbf{v}}_{1}^{\textrm{SRP}},\dots,\hat{\mathbf{v}}_{S}^{\textrm{SRP}}$ are estimated as those corresponding to the $S$ highest SRP-PHAT values on the finer grids.
For position estimation, the functional in (64) was first evaluated on a coarse grid with a 10 cm resolution along the x-, y-, and z-axes (i.e., $59\times 59\times 23=80,063$ grid points). The $\beta=3$ coarse grid points with the highest SRP-PHAT values were then re-evaluated on a finer grid with a 1 cm resolution around those points, within a 20 cm $\times$ 20 cm $\times$ 20 cm cube (i.e., $3\times 21^{3}=27,783$ grid points). After estimating the first source position $\hat{\mathbf{p}}_{1}^{\textrm{SRP}}$ corresponding to the largest value of the fine grids, for the second source only coarse grid points at least 50 cm from the coarse estimate corresponding to the first source position were considered to avoid considering high SRP-PHAT values in the neighbourhood of $\hat{\mathbf{p}}_{1}^{\textrm{SRP}}$ as separate sources (which could sometimes result in additional grid points being considered). For DOA estimation, the functional in (65) was first evaluated on a coarse grid with a $5^{\circ}$ resolution in azimuth ($\theta$) and elevation ($\psi$) (i.e., $(72\times 35)+2=2,522$ grid points - noting that for elevation angles 0 and 180 the DOA is independent of the azimuth angle). The $\beta=2$ coarse grid points with the highest SRP-PHAT values were then re-evaluated on a finer grid with a $0.5^{\circ}$ resolution, within $10^{\circ}$ in azimuth and elevation angles (i.e., $2\times 21^{2}=882$ grid points). After estimating the azimuth $\hat{\theta}_{1}^{\textrm{SRP}}$ and elevation $\hat{\psi}_{1}^{\textrm{SRP}}$ of the first source, for the second source only coarse grid points at least $20^{\circ}$ away from the coarse estimate corresponding to the first source DOA were considered (sometimes resulting in an additional coarse grid point being considered).
VI-D Experiment 1: Position Estimation
This section compares the position estimation accuracy of the proposed EDM-based method and the baseline SRP-based method for the acoustic scenarios with spatially distributed microphones described in Section VI-A. To evaluate the performance, we consider the position estimation error for each source, defined as
[…]
where the estimated source positions $\hat{\mathbf{p}}_{1}$ or $\hat{\mathbf{p}}_{2}$ are assigned to true source positions ${\mathbf{p}}_{1}$ or ${\mathbf{p}}_{2}$ using greedy assignment [47] (i.e., assigning each estimated source to the closest true source sequentially, in order of increasing $\varepsilon_{s}^{\textrm{pos}}$). Fig. 5 presents box plots of the position estimation errors $\varepsilon_{1}^{\textrm{pos}}$ and $\varepsilon_{2}^{\textrm{pos}}$ over 100 simulated scenarios, considering different distances $d_{\textrm{c}1}$ between source 1 and the array centroid.
Figure 5: Box plots of position estimation errors for both sources. The position estimation errors are shown for the EDM-based and SRP-based methods, for different distances $d_{\textrm{c}1}$ between source 1 and the array centroid ($d_{\textrm{c}2}=2$ m, 2 m cube). The red numbers at the top denote the number of results outside of the plotted range.
Table I summarizes the median position estimation errors for both sources.
TABLE I: Median position estimation errors for the EDM-based and SRP-based methods for different distances $d_{\textrm{c}1}$ between source 1 and the array centroid.
Median position estimation error [cm]
$d_{\textrm{c}1}[m]$ EDM, $\overline{\varepsilon}_{1}^{\textrm{pos}}$ EDM, $\overline{\varepsilon}_{2}^{\textrm{pos}}$ SRP, $\overline{\varepsilon}_{1}^{\textrm{pos}}$ SRP, $\overline{\varepsilon}_{2}^{\textrm{pos}}$
0 0.1 0.8 58.0 197.2
11 0.2 0.7 4.1 136.1
22 0.6 0.7 6.4 7.7
33 2.3 0.6 200.0 3.7
44 9.4 0.6 350.6 3.1
As can be observed from Table I, for all source distances $d_{\textrm{c}1}$ the proposed EDM-based method yields considerably lower median position estimation errors for both sources than the SRP-based method. In addition, it can be observed from Fig. 5 that the spread of the box plots is much smaller for the EDM-based method than for the SRP-based method for all source configurations. Furthermore, it is interesting to note that both methods behave quite differently for different source configurations. As can be observed from Table I, the median position estimation error for the EDM-based method mainly depends on source distance, i.e., increases for source 1 with increasing $d_{\textrm{c}1}$, whereas it remains relatively constant for source 2 (at $d_{\textrm{c}2}=2$ m). On the other hand, the median position estimation error for the SRP-based method depends both on the absolute distances between the sources and the array centroid as well as on the relative distance between the sources. For instance, it can be clearly observed that the median position estimation error is larger for the source that is farther away from the array centroid (e.g., for $d_{\textrm{c}1}=3$ m, $\overline{\varepsilon}_{1}^{\textrm{pos}}(\mathbf{s})=200$ cm and $\overline{\varepsilon}_{2}^{\textrm{pos}}(\mathbf{s})=3.7$ cm). This can be explained by the peaks of the farther source in the SRP functional being overpowered by the peaks of the closer source. In addition, when both sources are close to the microphones, the median position estimation error for both sources is large (e.g., for $d_{\textrm{c}1}=0$ m, $\overline{\varepsilon}_{1}^{\textrm{pos}}(\mathbf{s})=58$ cm and $\overline{\varepsilon}_{2}^{\textrm{pos}}(\mathbf{s})=197.2$ cm). This corresponds to [21, 24], where it was shown that the accuracy of SRP-based position estimation degrades for sources located close to the microphone array because the peaks in the SRP-PHAT functional become narrow. This would necessitate a high 3D grid resolution to detect these peaks, which is computationally impractical in most applications. In contrast, the EDM-based method only requires a one-dimensional grid search on the distance variable $\alpha$, which is computationally more efficient to optimize at a high resolution.
The average run times of the EDM-based and SRP-based position estimation methods are shown in Table II, when processing 5 second signals on a Ryzen 5900x processor. The EDM-based method is about 380 times faster than the SRP-based method, which can be attributed to the vast difference in search-space between both methods. Rather than jointly optimizing a functional in three position variables using all microphone pairs, the EDM-based method optimizes a cost function in a single continuous variable, based on estimated TDOAs between the reference microphone and the other microphones.
TABLE II: Average run times of the EDM-based and SRP-based source position and DOA estimation methods.
Method run time [s]
EDM SRP
Two-source position estimation 1.6 609.3
Two-source DOA estimation 1.1 27.2
In conclusion, the consistently small median errors and small error distributions demonstrate the versatility of the proposed EDM-based multi-source position estimation method across a wide range of configurations and with low computational cost.
VI-E Experiment 2: DOA Estimation
This section compares the DOA estimation accuracy of the proposed EDM-based method and the baseline SRP-based method for the acoustic scenarios with compact microphone arrays described in Section VI-A. To evaluate the performance, we consider the DOA estimation error for each source, defined as
[…]
where the estimated DOA vectors $\hat{\mathbf{v}}_{1}$ or $\hat{\mathbf{v}}_{2}$ are assigned to true DOA vectors ${\mathbf{v}}_{1}$ or ${\mathbf{v}}_{2}$ using greedy assignment. Fig. 6 presents box plots of the DOA estimation errors $\varepsilon_{1}^{\textrm{DOA}}$ and $\varepsilon_{2}^{\textrm{DOA}}$ over 100 simulated scenarios, considering different distances $d_{\textrm{c}1}$ between source 1 and the array centroid.
Figure 6: Box plots of DOA estimation errors for both sources. The DOA estimation errors are shown for the EDM-based and SRP-based methods, for different distances $d_{\textrm{c}1}$ between source 1 and the array centroid ($d_{\textrm{c}2}=2$ m, 10 cm cube). The red numbers at the top denote the number of results outside of the plotted range.
Table III summarizes the median DOA estimation errors for both sources.
TABLE III: Median DOA estimation errors for the EDM-based and SRP-based methods for different distances $d_{\textrm{c}1}$ between source 1 and the array centroid.
Median DOA estimation error []
$d_{\textrm{c}1}[m]$ EDM, $\overline{\varepsilon}_{1}^{\textrm{DOA}}$ EDM, $\overline{\varepsilon}_{2}^{\textrm{DOA}}$ SRP, $\overline{\varepsilon}_{1}^{\textrm{DOA}}$ SRP, $\overline{\varepsilon}_{2}^{\textrm{DOA}}$
$0.5$ 1.2 2.9 1.3 51.1
11 0.8 1.8 1.2 22.8
22 1.1 1.0 2.4 2.5
33 2.3 0.9 4.5 1.8
44 3.6 0.9 11.7 1.7
As can be observed from Table III for all source distances $d_{\textrm{c}1}$, the proposed EDM-based method yields median DOA estimation errors below $4^{\circ}$ for both sources, which are consistently smaller than the median DOA estimation errors obtained by the SRP-PHAT method. In addition, it can be observed from Fig. 6 that the spread of the box plots is smaller for the EDM-based method than for the SRP-based method for all source configurations. It is also interesting to note that for both methods the median DOA estimation error is larger for the source that is farther away from the array centroid. This can be explained by the fact that the amplitudes of the peaks in the GCC-PHAT and SRP-PHAT functionals corresponding to both sources depend on their signal power ratio. This effect is much more pronounced for the SRP-based method than for the EDM-based method. For instance, when source 1 is close to the microphone array ($d_{\textrm{c}1}=0.5$ m), the median DOA estimation error for source 1 is smaller than the median DOA estimation error for source 2 ($1.2^{\circ}$ and $2.9^{\circ}$ for the EDM-based method; $1.3^{\circ}$ and $51.1^{\circ}$ for the SRP-based method). In contrast, when source 1 is far from the microphone array ($d_{\textrm{c}1}=4$ m), the median DOA estimation error for source 1 is larger than the median DOA estimation error for source 2 ($3.6^{\circ}$ and $0.9^{\circ}$ for the EDM-based method; $11.7^{\circ}$ and $1.7^{\circ}$ for the SRP-based method). When both sources are equi-distant ($d_{\textrm{c}1}=d_{\textrm{c}2}=2$ m), both methods achieve similar median DOA estimation errors for both sources ($1.1^{\circ}$ and $1.0^{\circ}$ for the EDM-based method; $2.4^{\circ}$ and $2.5^{\circ}$ for the SRP-based method). For $d_{\textrm{c}1}=0.5$ m, it can be seen that the median DOA estimation errors are also slightly larger than for $d_{\textrm{c}1}=1$ m, reflecting that the far field model assumption becomes less valid for small distances between the source and the microphones.
The average run times of the EDM-based and SRP-based DOA estimation are shown in Table II. The EDM-based DOA method is about 25 times faster than the SRP-based method, which can again be attributed to the large difference in search-space between both methods. Rather than jointly optimizing a functional in two angle variables using all microphone pairs, the EDM-based method completely eliminates the need for continuous variable optimization, using a cost function based on estimated TDOAs between the reference microphone and the other microphones.
Similarly as for the position estimation results, the consistently small median DOA estimation errors and small error distributions demonstrate the versatility of the proposed EDM-based multi-source DOA estimation method across a wide range of source configurations, and with low computational cost.
VII Conclusions
In this paper, we have proposed novel TDOA-based multi-source position and DOA estimation methods that exploit properties of Euclidean distance matrices. For 3D position estimation, the proposed method requires optimizing only a single continuous variable instead of jointly optimizing three position variables as in SRP-based methods. This variable, representing the distance between each source and a reference microphone, is optimized for each combination of candidate TDOA estimates with a cost function based on the eigenvalues of the Gram matrix associated with the EDM. The estimated relative source positions, obtained from the Gram matrices corresponding to the optimal cost function values, are then mapped to absolute source positions by solving an orthogonal Procrustes problem for each source. For 3D DOA estimation, we define a relative coordinate system for each source, with one axis aligned to the DOA vector. The optimal set of candidate TDOA estimates is determined by minimizing a cost function based on the eigenvalues of a rank-reduced Gram matrix, requiring no continuous variable optimization. The source DOA vectors are then estimated by mapping the relative microphone positions, obtained from the rank-reduced Gram matrices, to the absolute microphone positions, for each source. For two sources in a noisy and reverberant environment, experimental results across a wide range of source and microphone configurations, including compact as well as distributed microphone arrays, show that the proposed EDM-based methods consistently outperform the SRP-based methods in terms of position and DOA estimation accuracy. Furthermore, due to the reduction in continuous variables to be optimized, the computational cost of the EDM-based methods is considerably lower than the SRP-based methods.